The following table gives the map coordinates and the shipping loads for a set of cities that we wish to connect through a central hub.                                                                                                 City Map Coordinates ​(​x,y​) Shipping Load Upper A ​(6​,11​) 5 Upper B ​(5​,8​) 10 Upper C ​(4​,8​) 14 Upper D ​(7​,5​) 6 Upper E ​(8​,8​) 14 Upper F ​(3​,1​) 11 Upper G ​(3​,6​) 4 ​a) For the location of the proposed new central​ hub, the coordinates should be​ near: x​ = 5.23 ​(round your response to two decimal​ places). y​ = 6.64 ​(round your response to two decimal​ places). ​b) If the shipments from city A double​, for the location of the proposed new central​ hub, the coordinates should be​ near: x​ = nothing ​(round your response to two decimal​ places).
Answer #1

(a)

City - j Xj Yj Lj Xj.Lj Yj.Lj
A 6 11 5 30 55
B 5 8 10 50 80
C 4 8 14 56 112
D 7 5 6 42 30
E 8 8 14 112 112
F 3 1 11 33 11
G 3 6 4 12 24
Totals 64 335 424
Centroid
Cx 335/64 = 5.23
Cy 424/64 = 6.63

(b)

City - j Xj Yj Lj Xj.Lj Yj.Lj
A 6 11 10 60 110
B 5 8 10 50 80
C 4 8 14 56 112
D 7 5 6 42 30
E 8 8 14 112 112
F 3 1 11 33 11
G 3 6 4 12 24
Totals 69 365 479
Centroid
Cx 365/69 = 5.29
Cy 479/69 = 6.94